OpenStudy (anonymous):

Solve the equation on the interval 0,2pi 2sin^2x=3(1-cosx)

OpenStudy (shubhamsrg):

you have this : 2(1- cos^2 x) = 3 -3cos x its a quad eqn in cos x.. just solve for cos x,,and then for x..

OpenStudy (shubhamsrg):

altough you can directly find out cos x via hit and trial..ain't that difficult,,try it :D (one of those traditional values)

OpenStudy (shubhamsrg):

you may use this: 1 - cos^2 x = (1+ cos x)(1- cos x) note that you cant directly cancel (1-cos x) as it equals to 0 (which you'll find out later)

OpenStudy (anonymous):

how did you get the (1+cosx)(1-cosx)?

OpenStudy (shubhamsrg):

a^2 - b^2 = (a+b)(a-b) <--- clearer now buddy?

OpenStudy (anonymous):

ohhh i see I thought you turned the 3-3cosx into that but thats in place of the 1-cos^2x

OpenStudy (shubhamsrg):

i didnt get your query..?? come again..

OpenStudy (shubhamsrg):

i just wrote (1)^2 - (cos x)^2 as (1+cos x)(1- cos x) that shouldnt be much of a problem..

OpenStudy (anonymous):

Yeah I see what you mean now. How would I solve for x now?

OpenStudy (shubhamsrg):

what values of cos x did you get ?

OpenStudy (anonymous):

none yet I'm stuck at that part

OpenStudy (shubhamsrg):

which part..write down where did you reach..??

OpenStudy (anonymous):

I've only gotten to 2(1+cosx)(1-cosx)=3(1-cosx) you said I can't cancel right

OpenStudy (shubhamsrg):

yes you cant,,hmm bring all terms on 1 side,,lets say LHS..now take (1-cos x) common,,what do you get?

OpenStudy (shubhamsrg):

??

OpenStudy (anonymous):

2/3(1+cosx)(1-cosx)(1-cosx)?

OpenStudy (shubhamsrg):

not dividing,,i meant subtracting :D subtract 3(1-cosx) from both sides,,and then take (1-cos x) common

OpenStudy (shubhamsrg):

<ding> ? comeon buddy,,it isnt that tought..make an attempt..go for it..

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

Would it just be 2(1-cos^2x)-3(1-cosx)?

OpenStudy (shubhamsrg):

yes ,,thats right,,now expand (1- cos^2 x) as (1-cos x)(1+cos x)..take out (1-cos x) common

OpenStudy (anonymous):

ok so 2(1+cosx)-3=0?

OpenStudy (shubhamsrg):

actually since you cant cancel (1-cos x),,you'll have this : (1- cos x)( 2 (1+cosx) -3) =0 so you get : 1-cosx =0 and 2(1+cos x) -3 =0 what values of cos x you get from here?

OpenStudy (anonymous):

cos x = -1 cos x = 1/2?

OpenStudy (shubhamsrg):

cos x=1/2 is right,, but -1 ?? it should be cos x=1..check again..

OpenStudy (anonymous):

oh youre right I forgot to include the negative sign in front of cos which changes them both to positive

OpenStudy (anonymous):

so then my final answers would be pi/3 and 5pi/3, since 2pi cant be included?

OpenStudy (shubhamsrg):

yep you are right,,since 0 and 2pi are not included,,these would be the ans.

OpenStudy (anonymous):

awesome. Thanks a lot for being patient im so tired so it took me a while lol

OpenStudy (shubhamsrg):

nevermind,,glad to help ^_^